OpenStudy (anonymous):

What is the equation of the line passing through (–3, –9) and parallel to the line y = x + 6 in standard form? plss help

hero (hero):

the equation will have the same slope, therefore use \[y-y_1 = m(x-x_1) \] Replace m with 1 and replace x1 and y1 with the given point into the equation, then simplify.

hero (hero):

Let me know what you get

OpenStudy (anonymous):

uhhh idkk im so confusedddd :(

hero (hero):

y - (-9) = 1(x - (-3))

hero (hero):

Are you still confused?

OpenStudy (anonymous):

um kindaa

hero (hero):

What exactly are you confused about?

OpenStudy (anonymous):

wait would it be y+9=x+3?

hero (hero):

yes...then just put it in terms of y

OpenStudy (anonymous):

what do u mean?

hero (hero):

When they say "put an equation in terms of y", it simply means to "solve for y".

OpenStudy (anonymous):

oh so y=x-6?

OpenStudy (anonymous):

but thats not any of the options x + y = –6 x – y = –6 x – y = 6 x + y = 6 see

OpenStudy (anonymous):

?

hero (hero):

well, what you do is, just simply subtract 6 from both sides, then subtract x from both sides....see what you get afterwards...

hero (hero):

This is kind of tricky....you'll have to do a bit of algebra to get it right

hero (hero):

You started out with y +9 = x + 3 if you subtract 9 from both sides, you get y = x -6 If you subtract x from both sides, you get y - x = -6 Now, switch the x and y: -x + y = -6 Now factor -1 from -x + y: -(x -y) = -6 Now divide both sides by -1: x - y = 6 So the final answer is: x - y = 6

OpenStudy (anonymous):

idk.